1/3

Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,

Der Werwolf The Werewolf The Banshee (An Approach)

Le loup-garou

tradukita de Alexander Gross  tradukita de Max Knight  tradukita de R. Platteau

Ein Werwolf eines Nachts ent-

wich von Weib und Kind und sich

begab an eines Dorfschullehrers Grab und bat ihn: �Bitte, beuge

mich!“ A Werewolf, troubled by his name, Left wife and brood one night and ca-

me To a hidden graveyard to enlist The aid of a long-dead philologist.

One night, a banshee slunk away from mate and child, and in the

gloom went to a village teacher’s tomb, requesting him: “Inflect me,

pray.”

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître d’école Et lui dit, “ sur moi dis tout, c’est ton rôle

! ”

Der Dorfschulmeister stieg hin-

auf auf seines Blechschilds Mes-

singknauf und sprach zum Wolf, der seine

Pfoten geduldig kreuzte vor dem To-

ten:

“Oh sage, wake up, please don’t berate

me,” He howled sadly, “Just conjugate me.” The seer arose a bit unsteady Yawned twice, wheezed once, and then

was ready.

The village teacher climbed up

straight upon his grave stone with its pla-

te and to the apparition said who meekly knelt before the

dead:

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître d’école Et lui dit, “ sur moi dis tout, c’est ton rôle

! ”

�Der Werwolf“ – sprach der gu-

te Mann, �des Werwolfs, Genitiv sodann, dem Werwolf, Dativ, wie man’s

nennt, den Werwolf, - damit hat’s ein

End!“ “Well, ‘Werewolf’ is your plural past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present ten-

se, And ‘Iswolf,’ ‘Arewolf’ in this same

sense.”

“The banSHEE, in the subject’s

place; the banHERS, the possessive ca-

se. The banHER, next, is what they

call objective case–and that is all.”

“ - On peut à ton sujet puisque nous y som-

mes Se demander encor bien des choses en som-

me. Loup-garoù ? mais avant Il faut se dire : loup-garquand ?

…  …  …  …

2/3

Dem Werwolf schmeichelten

die Fälle er rollte seine Augenbälle. �Indessen“, bat er, �füge doch zur Einzahl auch die Mehrzahl

noch!“ “I know that–I’m no mental cripple– The future form and participle Are what I crave,” the beast replied. The scholar paused–again he tried:

The banshee marveled at the ca-

ses and writhed with pleasure, ma-

king faces, but said: “You did not add, so far, the plural to the singular!”

Loup-garcomment ? et puis

Bien évidemment loup-garqui ? ”

Toutes ces questions semblèrent au loup bi-

en absconses, Car en fait ce qu’il voulait c’était les

réponses.

Der Dorfschullehrer aber mus-

ste gestehn, dass er von ihr nichts

wusste. Zwar Wölfe gab’s in großer

Schar, doch �Wer“ gab’s nur im Sinu-

lar.

“A ‘Will-be-wolf?’ It’s just too long: ‘Shall-be-wolf?’ ‘Has-been-wolf?’ Ut-

terly wrong! Such words are wounds beyond all

suture– I’m sorry, but you have no future.”

The teacher, though, admitted

then that this was not within his ken. “While bans are frequent”, he ad-

vised, “A she cannot be plurized.”

“ - Point de réponses, dit le puits de savoir, Etre imaginaire, nul ne peut te voir ! ” A ces mots l’animal éclata tout en larmes. “ - Oh!, même les baisers de ma douce moi-

tié

Der Wolf erhob sich

tränenblind – er hatte ja doch Weib und

Kind! Doch da er keine Gelehrter

eben, so schied er dankend und erge-

ben.

The Werewolf knew better–his sons

still slept At home, and homewards now he

crept, Happy, humble, without apology For such folly of philology.

The banshee, rising clammily, wailed: “What about my family?” Then, being not a learned crea-

ture, said humbly “Thanks” and left

the teacher.

Ne pourrons plus me faire me sentir tout

entier. Rester un loup-garou est un sort qui

m’alarme. ” Mais comme le loup était bien poli Il remercia et puis repartit.

…  …  …  …

3/3

Verkinto de tiu ĉi Germana poemo estas Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Francan de R. Platteau.

Arg-825-1655 (2013-02-13 21:56:38)

Tiu ĉi francigo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .